# neurips_data_2024.bib

### v1

@@ Appendix B Datasets
  • Enron [shetty2004enron] is a bipartite continuous-time graph where nodes are users and the temporal edges represent emails sent between users. Emails with multiple recipients are recorded as separate and simultaneously occurring edges, one per recipient. The temporal edges are resolved at the second level and the dataset spans approximately 3.6 years.
  • UCI [uci-dataset] is a unipartite continuous-time social network dataset from an online platform at the University of California at Irvine. The nodes represent students and the timestamped edges represent communication between the students. The dataset spans approximately six and a half months.
  • MOOC (massive open online course) [PredictingDynamicEmbedding2019kumar] is a bipartite continuous-time graph where nodes represent users and units in an online course, such as problems or videos. Temporal edges are resolved at the second level and encode when a user interacts with a unit of the online course. The dataset spans approximately one month.
  • Wikipedia (Wiki.) [PredictingDynamicEmbedding2019kumar] is a bipartite continuous-time graph where nodes represent editors and Wikipedia articles. The timestamped edges are resolved at the second level and represent when an editor has edited an article. The dataset spans approximately one month.
[…]
  • Myket [EffectChoosingLoss2023loghmania] is a bipartite continuous-time graph where nodes represent user and Android applications. The timestamped edges represent when a user installed an application. The dataset spans approximately six and a half months.
  • Social Evolution (Social) [social-evo] is a unipartite continuous-time graph of the proximity between the students in a dormitory, collected between October 2008 and May 2009 using mobile phones. Temporal edges connect students when they are in proximity and are resolved at the second level.
  • Reddit [PredictingDynamicEmbedding2019kumar] is a bipartite continuous-time graph where nodes represent Reddit users and their posts. The timestamped edges are resolved at the second level and represent when a user has made a post on Reddit. The dataset spans approximately one month.
  • UN Vote (UN V.)[un-voting-dataset, BetterEvaluationDynamic2022poursafaei] is a weighted unipartite discrete-time graph of votes in the United Nations General Assembly between 1946 and 2020. Nodes represent countries and edges connect countries if they both vote “yes”. The dataset is resolved at the year level and edge weights represent how many times the two connected countries have both voted “yes” in the same vote.
  • US Legislators (US L.) [10.1145/3394486.3403077, FOWLER2006454, BetterEvaluationDynamic2022poursafaei] is a weighted unipartite discrete-time graph of interactions between legislators in the US Senate. Nodes represent legislators and edges represent co-sponsorship, i.e., edges connect legislators who co-sponsor the same bill. The dataset is resolved at the congress level and edge weights encode the number of co-sponsorships during a congress.
  • UN Trade (UN Tr.) [10.1093/biosci/biu225, BetterEvaluationDynamic2022poursafaei] is a directed and weighted unipartite discrete-time graph of food and agricultural trade between countries where nodes represent countries. The dataset spans 30 years and is resolved at the year level. Weighted edges encode the sum of normalized agriculture imports or exports between two countries during a given year.
  • Canadian Parliament (Can. P.) [10.1145/3394486.3403077, BetterEvaluationDynamic2022poursafaei] is a weighted unipartite discrete-time political network where nodes represent Members of the Canadian Parliament (MPs) and an edge between two MPs means that they have both voted “yes” on a bill. The dataset is resolved at the year level and the edges’ weights represent how often the two connected MPs voted “yes” on the same bill during a year.
  • Flights [flights-dataset, BetterEvaluationDynamic2022poursafaei] is a directed and weighted unipartite discrete-time graph where nodes represent airports and edges represent flights during the COVID-19 pandemic. The edges are resolved at the day level and their weights are given by the number of flights between two airports during the respective day.
  • Contacts (Cont.) [Sapiezynski2019, BetterEvaluationDynamic2022poursafaei] is a unipartite discrete-time proximity network between university students. Nodes represent students who are connected by an edge if they were in close proximity during a time window. The dataset is resolved at the 5-minute level and spans one month.
Appendix C Experimental details

### v2

@@ C.4 Number of Links per Timestamp
C.5 Window Sizes for Different Window Durations
Figure 8 shows the distribution of the number of links per time window for continuous- and discrete-time data sets. We can see that the distribution of window sizes depends on the distribution of edges across time as visualised in Figure 5. Datasets with a more uniform activity pattern such as Wikipedia and Reddit show a similar number of edges across windows of the same duration. Datasets with inhomogeneously distributed activities, e.g. Enron or UCI, exhibit a wide range of window sizes. For large values of $h$, i.e. horizons that include more than half of the observed time, the window sizes diverge because the window of the last time window that still includes observed links gets smaller.
(a) The number of links per window in continuous-time temporal graphs for horizons ranging from one second to one year. The $x$-axis is labelled in hours.
(b) Window sizes for discrete-time temporal graphs where $h$ represents the number of snapshots.
Figure 8: Window sizes, i.e. number of links per time window, for different horizons.
Appendix D Link Prediction using Different Batch Sizes
The results in Table 5 and 6 show the differences in AUC-ROC and average precision scores using two different batch sizes. While the same models with the same trained weights are used for both evaluations, the results show substantial differences in performance. This is in line with our expectations as described in Problem (4).
As expected due to Problem (1), the performance differences are especially large for memory-based models and the performance drops almost consistently for all datasets when using a larger batch size. Other models show no clear preference for whether larger or smaller batch sizes lead to better performance. This drop in performance for memory-based models specifically can be explained because they update their internal node representations after processing each batch, thus, larger batches lead to less frequent updates and more “out-of-date” internal representations (Su et al., 2024; Feldman & Baskin, 2024). In other words, larger batches lead to a larger information loss because more edges at the beginning of each batch are missing in the internal representation for the prediction of edges at the end of each batch.
Table 5: Mean AUC-ROC performance for dynamic link prediction ($b=800$) over five runs for the continuous- (top) and discrete-time (bottom) datasets. Values in parenthesis show the relative change compared to an evaluation using batch size $b=200$.
Enron 74.8( $\downarrow$3.4%) 67.9( $\downarrow$7.7%) 68.4( $\uparrow$0.5%) 59.4( $\uparrow$1.2%) 67.1( $\uparrow$1.1%) 78.5( $\downarrow$1.7%) 68.4( $\uparrow$1.1%) 82.2( $\uparrow$1.1%) 77.5( $\uparrow$1.4%) 2.1%±2.2%
UCI 79.9( $\downarrow$4.1%) 47.7( $\downarrow$7.3%) 61.0( $\downarrow$3.3%) 59.2( $\downarrow$0.7%) 57.6( $\downarrow$0.9%) 64.6( $\downarrow$6.6%) 59.6( $\downarrow$0.6%) 79.9( $\downarrow$0.9%) 76.0( $\downarrow$0.2%) 2.7%±2.7%
MOOC 82.0( $\downarrow$3.4%) 77.9( $\downarrow$3.5%) 87.5( $\downarrow$1.1%) 82.5( $\uparrow$0.2%) 70.5( $\uparrow$0.1%) 57.2( $\downarrow$7.6%) 72.6( $\downarrow$0.0%) 74.5( $\uparrow$0.1%) 81.3( $\uparrow$0.1%) 1.8%±2.6%
Wiki. 77.4( $\downarrow$5.4%) 74.5( $\downarrow$4.9%) 83.7( $\downarrow$0.5%) 83.5( $\uparrow$0.1%) 72.0( $\uparrow$0.5%) 76.5( $\downarrow$0.8%) 85.2( $\uparrow$0.1%) 88.0( $\uparrow$0.2%) 80.4( $\uparrow$0.5%) 1.4%±2.1%
LastFM 76.5( $\downarrow$2.0%) 70.4( $\downarrow$1.8%) 80.1( $\downarrow$0.8%) 68.4( $\uparrow$0.0%) 67.9( $\downarrow$0.2%) 77.4( $\downarrow$1.0%) 64.3( $\uparrow$0.1%) 65.9( $\downarrow$0.1%) 78.9( $\downarrow$0.0%) 0.7%±0.8%
Myket 64.4( $\uparrow$0.7%) 63.6( $\downarrow$1.0%) 60.7( $\downarrow$0.7%) 57.4( $\downarrow$0.3%) 32.6( $\uparrow$0.3%) 52.0( $\uparrow$0.1%) 58.4( $\downarrow$0.0%) 59.4( $\downarrow$0.2%) 33.0( $\uparrow$0.5%) 0.4%±0.3%
Social 87.3( $\downarrow$4.4%) 88.2( $\downarrow$4.8%) 92.0( $\uparrow$0.3%) 92.7( $\uparrow$0.1%) 87.8( $\uparrow$0.2%) 86.0( $\uparrow$0.2%) 95.3( $\uparrow$0.1%) 94.1( $\uparrow$0.1%) 97.4( $\uparrow$0.1%) 1.1%±2.0%
Reddit 79.0( $\downarrow$1.9%) 76.4( $\downarrow$3.9%) 80.3( $\downarrow$0.1%) 78.7( $\uparrow$0.1%) 80.3( $\uparrow$0.1%) 78.5( $\downarrow$0.2%) 76.2( $\uparrow$0.0%) 77.2( $\uparrow$0.1%) 80.2( $\uparrow$0.0%) 0.7%±1.3%
UN V. 71.3( $\downarrow$3.1%) 66.9( $\downarrow$7.8%) 67.2( $\downarrow$4.4%) 52.8( $\uparrow$0.0%) 50.2( $\uparrow$0.2%) 89.6( $\uparrow$0.1%) 53.0( $\downarrow$0.1%) 56.5( $\uparrow$0.4%) 63.3( $\uparrow$0.4%) 1.8%±2.7%
US L. 56.6( $\uparrow$0.6%) 73.5( $\downarrow$7.9%) 76.7( $\downarrow$8.7%) 78.6( $\uparrow$0.1%) 82.6( $\uparrow$0.9%) 67.2( $\downarrow$0.4%) 77.0( $\uparrow$1.9%) 90.3( $\uparrow$0.1%) 90.2( $\uparrow$0.8%) 2.4%±3.4%
UN Tr. 65.7( $\downarrow$0.6%) 63.2( $\uparrow$0.1%) 61.5( $\downarrow$2.6%) 61.4( $\downarrow$0.5%) 64.6( $\downarrow$0.2%) 86.5( $\uparrow$0.1%) 60.9( $\downarrow$0.1%) 66.2( $\downarrow$0.1%) 68.2( $\downarrow$0.2%) 0.5%±0.8%
Can. P. 62.2( $\downarrow$2.7%) 64.6( $\downarrow$3.0%) 72.9( $\downarrow$0.7%) 71.1( $\downarrow$0.7%) 67.9( $\downarrow$0.2%) 62.1( $\downarrow$1.4%) 68.2( $\uparrow$0.0%) 81.0( $\downarrow$0.2%) 93.3( $\downarrow$4.5%) 1.5%±1.6%
Flights 68.9( $\downarrow$0.9%) 68.4( $\downarrow$0.9%) 68.5( $\downarrow$0.4%) 72.7( $\uparrow$0.1%) 66.0( $\uparrow$0.0%) 74.6( $\uparrow$0.0%) 70.6( $\downarrow$0.0%) 70.7( $\downarrow$0.0%) 68.7( $\downarrow$0.0%) 0.3%±0.4%
Cont. 93.4( $\downarrow$2.2%) 93.3( $\downarrow$2.2%) 95.5( $\downarrow$0.6%) 95.5( $\uparrow$0.1%) 83.3( $\uparrow$0.1%) 91.7( $\downarrow$0.5%) 94.6( $\uparrow$0.1%) 94.2( $\uparrow$0.1%) 97.2( $\uparrow$0.1%) 0.7%±0.9%
$\mu\pm\sigma$ 2.5%±1.5% 4.1%±2.8% 1.8%±2.4% 0.3%±0.4% 0.4%±0.4% 1.5%±2.4% 0.3%±0.6% 0.3%±0.3% 0.6%±1.2%
Table 6: Mean AP performance for dynamic link prediction ($b=800$) over five runs for the continuous- (top) and discrete-time (bottom) datasets. Values in parenthesis show the relative change compared to an evaluation using batch size $b=200$.
Enron 68.2( $\downarrow$5.4%) 62.4( $\downarrow$10.4%) 65.2( $\downarrow$0.2%) 63.2( $\uparrow$0.1%) 66.8( $\uparrow$1.2%) 75.8( $\downarrow$1.5%) 70.7( $\uparrow$0.6%) 83.1( $\uparrow$0.9%) 77.2( $\uparrow$1.0%) 2.4%±3.4%
UCI 77.5( $\downarrow$4.5%) 46.3( $\downarrow$5.5%) 70.1( $\downarrow$1.5%) 68.3( $\downarrow$0.4%) 65.2( $\uparrow$0.1%) 61.0( $\downarrow$6.2%) 68.9( $\downarrow$0.2%) 85.5( $\downarrow$0.4%) 81.1( $\uparrow$0.6%) 2.2%±2.5%
MOOC 80.8( $\downarrow$3.1%) 74.7( $\downarrow$3.1%) 85.6( $\downarrow$1.6%) 84.6( $\uparrow$0.1%) 73.5( $\uparrow$0.0%) 56.1( $\downarrow$7.5%) 78.3( $\downarrow$0.2%) 77.8( $\downarrow$0.1%) 82.4( $\uparrow$0.0%) 1.7%±2.5%
Wiki. 79.7( $\downarrow$5.3%) 77.3( $\downarrow$4.3%) 88.5( $\downarrow$0.3%) 87.9( $\uparrow$0.1%) 75.4( $\uparrow$0.6%) 72.6( $\downarrow$0.7%) 89.6( $\uparrow$0.1%) 91.3( $\uparrow$0.2%) 83.6( $\uparrow$0.7%) 1.4%±2.0%
LastFM 75.3( $\downarrow$3.0%) 68.6( $\downarrow$3.9%) 80.1( $\downarrow$0.3%) 75.2( $\uparrow$0.3%) 69.8( $\uparrow$0.0%) 72.5( $\downarrow$1.0%) 71.9( $\uparrow$0.3%) 74.2( $\uparrow$0.2%) 81.2( $\uparrow$0.2%) 1.0%±1.4%
Myket 63.7( $\uparrow$0.4%) 61.0( $\downarrow$2.4%) 61.6( $\downarrow$0.8%) 56.6( $\downarrow$0.9%) 45.0( $\downarrow$0.3%) 51.2( $\downarrow$0.1%) 57.9( $\downarrow$0.7%) 58.8( $\downarrow$0.6%) 44.6( $\downarrow$0.3%) 0.7%±0.7%
Social 83.2( $\downarrow$5.8%) 85.7( $\downarrow$6.5%) 93.5( $\uparrow$0.3%) 94.9( $\uparrow$0.1%) 86.3( $\uparrow$0.1%) 80.8( $\uparrow$0.3%) 96.3( $\uparrow$0.1%) 95.5( $\uparrow$0.0%) 97.5( $\uparrow$0.2%) 1.5%±2.7%
Reddit 78.3( $\downarrow$2.3%) 75.7( $\downarrow$4.4%) 80.3( $\downarrow$0.3%) 78.5( $\downarrow$0.2%) 81.1( $\downarrow$0.0%) 73.5( $\downarrow$0.2%) 76.5( $\downarrow$0.1%) 77.4( $\downarrow$0.1%) 82.7( $\downarrow$0.0%) 0.8%±1.5%
UN V. 65.1( $\downarrow$4.0%) 62.2( $\downarrow$8.3%) 63.1( $\downarrow$4.5%) 51.8( $\downarrow$0.8%) 50.3( $\downarrow$0.7%) 84.9( $\uparrow$0.1%) 52.9( $\downarrow$0.7%) 53.7( $\downarrow$0.5%) 60.0( $\uparrow$0.0%) 2.2%±2.8%
US L. 48.2( $\uparrow$0.0%) 67.8( $\downarrow$7.2%) 74.4( $\downarrow$8.3%) 70.8( $\downarrow$0.6%) 81.3( $\uparrow$0.7%) 63.1( $\downarrow$0.3%) 78.8( $\uparrow$1.8%) 86.2( $\uparrow$0.2%) 86.7( $\uparrow$1.1%) 2.2%±3.2%
UN Tr. 57.9( $\downarrow$1.7%) 58.5( $\downarrow$1.2%) 57.2( $\downarrow$2.8%) 57.2( $\downarrow$1.2%) 57.4( $\downarrow$0.8%) 81.1( $\uparrow$0.1%) 55.7( $\downarrow$0.9%) 63.7( $\downarrow$0.0%) 64.5( $\uparrow$0.4%) 1.0%±0.9%
Can. P. 51.4( $\downarrow$2.6%) 60.4( $\downarrow$2.2%) 69.1( $\uparrow$0.8%) 68.1( $\uparrow$0.6%) 64.5( $\uparrow$0.9%) 61.7( $\downarrow$3.3%) 65.0( $\uparrow$1.2%) 78.9( $\uparrow$2.2%) 93.1( $\downarrow$4.1%) 2.0%±1.2%
Flights 65.8( $\downarrow$1.4%) 65.7( $\downarrow$1.5%) 68.2( $\downarrow$0.2%) 73.0( $\uparrow$0.4%) 64.8( $\downarrow$0.3%) 70.4( $\downarrow$0.1%) 70.8( $\downarrow$0.0%) 71.4( $\uparrow$0.3%) 68.4( $\downarrow$0.2%) 0.5%±0.6%
Cont. 91.4( $\downarrow$3.0%) 93.3( $\downarrow$2.3%) 95.6( $\downarrow$0.6%) 96.1( $\uparrow$0.1%) 84.5( $\uparrow$0.1%) 88.4( $\downarrow$0.5%) 95.1( $\uparrow$0.1%) 94.4( $\uparrow$0.1%) 97.8( $\uparrow$0.1%) 0.8%±1.1%
$\mu\pm\sigma$ 3.0%±1.8% 4.5%±2.8% 1.6%±2.3% 0.4%±0.4% 0.4%±0.4% 1.6%±2.4% 0.5%±0.5% 0.4%±0.6% 0.6%±1.1%
Appendix E Evaluation based on Realistic Time Horizons
In our main experiments, we determined the time horizon based on the average number of links per time window to enable a fair comparison to the batch-based approach. In Section 3.1, we discussed that the time horizon for a realistic evaluation should instead be chosen carefully for each individual dataset. To provide an example of such a realistic evaluation, we will assign a reasonable horizon for each dataset used in this work. Note that this enables a fair model comparison since we use the same horizon for all models across each dataset. The temporal resolution for most discrete-time datasets is limited due to the data collection process. The US Legislators dataset for example contains one snapshot for each congress which provides a natural horizon of one snapshot. Thus, we will only consider continuous-time datasets and discrete-time datasets where the duration of a snapshot is not yet a natural time horizon for evaluation, i.e. Contacts.
We list the considered datasets and the chosen horizon with the reasoning behind it in the following:
  • Enron is a network of users with edges representing emails sent between them. We choose 24 hours as a reasonable horizon since 90% of all email replies are typically sent within a day (Kooti et al., 2015).
  • UCI is a social network based on student communications. We use 30 min as the horizon since users receiving a text message typically feel pressured to reply between the next 20 minutes and the end of the day (Aranda & Baig, 2018).
  • MOOC connects students to units of an online course based on their interactions. We set $h=6\text{ min}$ since Guo et al. (2014) recommend keeping the learning video of a unit shorter than this time frame.
  • Wikipedia represents the editing behaviour of users in a graph. Since editing Wikipedia articles is unpaid, we do not expect frequent interactions from each user. We assume that users come from different time zones and consider that the total duration of the dataset is only one month. Therefore, we see the typical working time of 8 hours as an appropriate time horizon to take into account that interactions won’t appear very frequently but there are still enough time windows for evaluation.
  • LastFM connects users to the songs they listen to. We select 24 hours as the horizon to evaluate this dataset on a task where the goal is to predict the songs that users will listen to tomorrow based on their listening behaviour during the last days.
  • Myket represents users and Android applications that are connected when an application is installed. Considering that, similar to the Wikipedia dataset, no frequent interactions are expected, we choose 24 hours as the time horizon since the total duration of the dataset is longer than the total duration of the Wikipedia dataset.
  • Social Evolution is a proximity network gathered from students in a dormitory. Thus, we select 2 hours – a typical duration of a lecture including breaks – as the time horizon.
  • Reddit contains the posting behaviour of Reddit users. Since dynamics in a social network are typically fast, we select a time horizon of 15 minutes.
  • Contacts: Similar to Social Evolution, we select 2 hours as the time horizon since the network also captures the proximity of students.
We use the trained models from the experiments of the main part of our work and reevaluate the specified datasets with the selected time horizons. The results are presented in Table 7 and 8. For completeness, both tables also include the performance scores of the discrete-time datasets that have not been reevaluated because we determined the horizon used above as realistic.
Table 7: Average AUC-ROC performance and standard deviation over five runs using the trained models from the above experiments and the window-based evaluation with the horizons specified above. The tables also include the datasets that have not been reevaluated using performance scores obtained from the main experiments’ time horizons.
Datasets JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
Enron 84.2 ± 4.8 80.4 ± 2.5 70.0 ± 4.5 68.3 ± 1.8 75.5 ± 0.4 83.1 ± 0.0 74.2 ± 4.8 87.9 ± 0.6 83.9 ± 0.6
UCI 89.4 ± 0.7 70.4 ± 2.8 64.0 ± 1.3 52.2 ± 1.1 56.0 ± 0.2 75.3 ± 0.0 55.3 ± 1.3 82.7 ± 0.9 75.4 ± 0.4
MOOC 84.6 ± 4.1 81.3 ± 3.3 87.6 ± 1.3 80.7 ± 0.9 69.3 ± 1.5 62.9 ± 0.0 68.1 ± 0.9 70.4 ± 1.6 79.9 ± 10.0
Wiki. 75.5 ± 0.8 73.2 ± 0.9 83.5 ± 0.6 83.5 ± 0.2 71.9 ± 0.8 76.1 ± 0.0 85.2 ± 0.6 87.9 ± 0.3 80.6 ± 1.6
LastFM 74.9 ± 2.1 67.5 ± 1.4 78.6 ± 2.9 66.0 ± 0.9 67.2 ± 0.3 77.5 ± 0.0 63.2 ± 6.7 60.6 ± 1.3 78.9 ± 0.6
Myket 64.5 ± 1.8 62.7 ± 3.1 60.4 ± 2.3 57.8 ± 0.4 32.2 ± 0.4 51.4 ± 0.0 58.3 ± 2.0 59.9 ± 0.3 32.8 ± 0.9
Social 87.0 ± 2.0 84.4 ± 4.7 92.3 ± 2.7 92.6 ± 0.5 86.7 ± 0.0 85.1 ± 0.0 94.7 ± 0.5 94.7 ± 0.2 97.4 ± 0.1
Avg. Rank 3.0 5.2 3.8 6.1 6.9 6.0 6.1 4.2 3.6
UN V. 54.0 ± 1.8 52.2 ± 2.0 51.3 ± 7.1 54.4 ± 3.6 53.7 ± 2.1 89.6 ± 0.0 53.4 ± 1.0 56.9 ± 1.6 65.2 ± 1.1
UN Tr. 57.7 ± 3.3 50.3 ± 1.4 54.3 ± 1.5 64.0 ± 1.3 67.6 ± 1.2 85.6 ± 0.0 63.7 ± 1.6 68.6 ± 2.6 70.7 ± 2.6
Flights 67.4 ± 2.0 66.0 ± 1.9 68.1 ± 1.7 72.6 ± 0.2 66.0 ± 1.7 74.6 ± 0.0 70.6 ± 0.1 70.7 ± 0.3 68.8 ± 1.0
Cont. 85.4 ± 0.5 74.5 ± 3.1 94.6 ± 0.6 96.0 ± 0.2 86.6 ± 0.1 85.8 ± 0.0 95.7 ± 0.4 95.2 ± 0.2 97.8 ± 0.0
Avg. Rank 7.3 8.0 6.5 3.5 5.5 4.2 5.2 2.7 2.2
The results using both the AUC-ROC as well as the average precision score mostly agree on a best-performing model, yet for different datasets, there is no clear winner among the models. For continuous-time datasets, JODIE, GraphMixer, and DyGFormer are among the best-performing models while EdgeBank, GraphMixer, and DyGFormer performed best for discrete-time data.
DyGFormer performs best for both Contacts and Social Evolution, suggesting that DyGFormer is best suited for proximity networks among all models. For UN Vote, UN Trade, and Flights, EdgeBank – a simple baseline model that predicts an edge if it has occurred before – is among the best. These datasets are highly repetitive because they are, e.g. based on a schedule or relations among countries that rarely change. Thus, since they are all outperformed by simple baselines, none of the proposed TGNN models adequately address the task of these datasets, i.e. finding edges that do not follow the schedule or some other reoccurring pattern. For other types of datasets like communication (e.g. Enron or UCI) or user-interaction networks (e.g. Wikipedia or MOOC), no clear patterns are visible.
Table 8: Mean average precision scores and standard deviation following Table 7.
Datasets JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
Enron 82.2 ± 4.8 80.0 ± 3.4 71.2 ± 4.2 72.6 ± 1.3 77.1 ± 0.3 81.6 ± 0.0 77.9 ± 2.6 89.2 ± 0.4 85.0 ± 0.6
UCI 93.0 ± 0.5 81.5 ± 1.9 78.0 ± 0.9 71.6 ± 0.7 73.3 ± 0.2 75.6 ± 0.0 73.2 ± 0.7 89.6 ± 0.5 85.2 ± 0.3
MOOC 84.1 ± 5.4 79.4 ± 3.8 87.4 ± 1.6 83.9 ± 0.8 73.8 ± 1.1 62.1 ± 0.0 75.8 ± 0.5 75.6 ± 0.9 82.6 ± 8.9
Wiki. 78.3 ± 1.2 76.4 ± 0.7 88.3 ± 0.4 88.0 ± 0.2 75.7 ± 1.1 72.4 ± 0.0 89.7 ± 0.4 91.3 ± 0.2 84.0 ± 1.2
LastFM 73.6 ± 2.0 64.8 ± 1.8 78.6 ± 3.7 73.0 ± 0.8 69.2 ± 0.5 72.9 ± 0.0 71.2 ± 6.7 71.0 ± 1.1 81.3 ± 0.9
Myket 62.9 ± 1.5 60.2 ± 1.9 61.1 ± 1.7 56.8 ± 0.4 44.9 ± 0.2 51.1 ± 0.0 57.8 ± 2.2 59.0 ± 0.2 44.4 ± 1.7
Social 82.5 ± 4.0 81.6 ± 5.7 94.0 ± 1.8 95.0 ± 0.3 85.8 ± 0.1 79.9 ± 0.0 96.2 ± 0.3 95.8 ± 0.2 97.8 ± 0.1
Avg. Rank 3.4 5.8 3.8 5.5 6.8 7.2 5.4 4.0 3.2
UN V. 52.6 ± 1.8 49.6 ± 1.9 49.7 ± 3.9 52.7 ± 2.6 52.4 ± 2.0 84.2 ± 0.0 52.4 ± 0.9 54.0 ± 1.4 62.4 ± 1.7
UN Tr. 52.7 ± 3.0 49.4 ± 0.9 53.2 ± 1.5 58.9 ± 2.7 59.2 ± 1.7 79.0 ± 0.0 57.5 ± 1.9 65.8 ± 1.9 67.1 ± 2.7
Flights 65.2 ± 2.7 63.9 ± 2.8 68.3 ± 2.2 73.5 ± 0.3 65.5 ± 1.3 70.4 ± 0.0 71.0 ± 0.4 71.9 ± 0.8 69.1 ± 1.8
Cont. 83.3 ± 0.7 69.7 ± 3.7 95.0 ± 0.8 96.9 ± 0.2 88.1 ± 0.2 82.3 ± 0.0 96.7 ± 0.5 95.8 ± 0.1 98.5 ± 0.0
Avg. Rank 7.7 8.3 6.3 3.3 5.2 4.7 4.8 2.5 2.2
Appendix F Experimental Details